OpenStudy (anonymous):

~I NEED THIS ASAP, PLEASE HELP ~ Paul plans to put concrete on a rectangular portion of his driveway. The portion is 12 feet long and 6 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Paul needs is $108.89. Which of the following is closest to the width of the portion of the driveway on which Paul plans to put concrete? [1 foot = 12inches; 1 yard = 3 feet]

OpenStudy (***[isuru]***):

Hi, 1 cubic yard = 27 cubit feet then we must find the no.of cubic yards .... for 1 cubic yard = $98 for "x" cubic yard = $108.89 1/x =98/108.89 x = 1.11112 Now we have to convert this in to cubic feet... 1 cubic yard = 27 cubic feet 1.11112 cubic yard = x cubic feet 1/1.11112 = 27/x x = 30.0024 cubic feet now width x height x length = 30.004 cubic feet w x 12 x (1/2) = 30.004 w = 30.004/6 = 5.0004 feet yep! there for the width of the yard is approximately 5 feet.. hope this will help ya !

OpenStudy (anonymous):

Oh my god, thank you.!!! <33

OpenStudy (***[isuru]***):

u r welcome !